Tagged order is 0.1 probability that there are less than three tagged order forms in the sample of four.
1
Expert's answer
2021-10-03T17:42:28-0400

Let "X=" the number of tagged order forms: "X\sim Bin(n, p)."

Given "n=4, p=0.1,q=1-p=1-0.1=0.9."


"P(X<3)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{4}{0}(0.1)^0(0.9)^{4-0}+\dbinom{4}{1}(0.1)^1(0.9)^{4-1}"

"+\dbinom{4}{2}(0.1)^2(0.9)^{4-2}=0.9963"